Write an equation in slope intercept form satifying the given conditions:
passes through (-4,7) and (2,4)

Answer by fractalier(2101)

Find the slope m first, then put it and one of the points into point-slope form and solve for y...here goes...
m = (y2 - y1) / (x2 - x1)
m = (4 - 7) / (2 - (-4))
m = -3/6
m = -1/2
Now
y - y1 = m(x - x1)
y - 4 = (-1/2)(x - 2)
y - 4 = (-1/2)x + 1
y = (-1/2)x + 5